# Exploring Time Delay Interferometry Ranging as a Practical Ranging Approach in the Bayesian Framework

### v3

@@ V Conclusions
In this paper, we explored the feasibility and performance of TDIR as a ranging technique independent of the PRN signal, considering the realistic scenarios including LFN, secondary noises, OB noise, clock noise, orbital motion, and clock desynchronization. MCMC sampling based on simulated data yielded a 5.28 ns ranging bias together with a 2 ns dispersion, lower than the threshold which is set according to the effect of LFN suppression.
Our research has shown promising prospects for the application of TDIR, which may serve as a cross-validation or backup for the traditional PRN-based approach. This approach is free of the imperfections of PRNR and the sophisticated PRNR processing steps. Additionally, the whole TDI data processing flow based on our TDIR algorithm has more relaxed requirement on the accuracy of clock synchronization.
Meanwhile, there are at least three issues that remain to be addressed in the future. Firstly, in our research, running the two-phase MCMC sampling took around $\mathcal{O}(5)$ days on a personal laptop, highlighting the imperative to enhance both the computational power and the efficiency of algorithm. Secondly, since TDIR is driven by the interferometric measurements, the components of data have a significant impact on the outcome. In the presence of GW signals, it is necessary to incorporate TDIR into the global fit pipeline, and the effect of joint estimation with GW source parameters necessitates further investigation. Thirdly, the majority of our research assumed six independent lasers, and a simple illustrative example was provided for the case with laser locking. More detailed discussion regarding the impacts and choice of different locking schemes also awaits future research.
@@ Appendix A Noise models
Appendix B An illustrative example for the case with laser locking
Figure 7: Similar to FIG. 5, but the green and orange corner plots represent the posteriors of the “no locking scheme” and “N1-LA12 locking scheme” cases, respectively. Both results are obtained based on LFN-only simulations.

### v4

@@ V Conclusions
In this paper, we explored the feasibility and performance of TDIR as a ranging approach independent of the PRN signal, considering the realistic scenarios including LFN, secondary noises, OB noise, clock noise, orbital motion, and clock desynchronization. MCMC sampling based on simulated data yielded a 5.28 ns ranging bias together with a 2 ns dispersion, lower than the threshold which is set according to the effect of LFN suppression.
Our research has shown promising prospects for the application of TDIR, which may serve as a cross-validation or backup for the traditional PRN code-based method. This approach is free of the imperfections of PRNR, and hence the sophisticated PRNR processing steps. Additionally, the whole TDI data processing flow based on our TDIR algorithm has more relaxed requirement on the accuracy of clock synchronization.
Meanwhile, there are at least four issues that remain to be addressed in the future. Firstly, in our research, running the two-phase MCMC sampling took $\sim\mathcal{O}(5)$ days on a personal laptop, highlighting the imperative to enhance both the computational power and the efficiency of algorithm. For example, using an optimization algorithm to perform a maximum likelihood estimation (MLE) may be a more efficient approach. After all, in practical applications, we are more concerned with the optimal values of ranges rather than their posterior distributions.
Secondly, since TDIR takes the interferometric data as the input, the compositions of data have significant impacts on the outcome. Although we have tried to avoid the encounter with GW signals by only using the high-frequency data, the possibility that there might still be signals can not be ruled out. In the presence of GW signals, it is necessary to incorporate TDIR into the global fit pipeline, and the effect of joint estimation with GW source parameters necessitates further investigation.
Thirdly, in the calculation of likelihood function, we have assumed that the noise models $P^{\rm 2nd}_{X}(f)$, as functions of frequency and armlengths, are known a prior, indicating that the secondary noises have been well calibrated. While, in practice, due to the variations of instruments and space environment, the noise models may evolve over time, thus TDIR may confront the situation of inaccurate or unknown noise models. In face of such situation, an iterative estimation may be a possible solution. To be specific, the 1st step is a MLE with $P^{\rm 2nd}_{X}(f)\equiv 1$, then we construct TDI combinations using the resulting TDIR, and subsequently evaluate the PSDs of these TDI combinations. These PSDs will serve as $P^{\rm 2nd}_{X}(f)$ and be input to the 2nd TDIR estimation step. As such, the iterative process of refining the estimates of TDIR and PSD will continue incrementally until convergence.
At last, the majority of our research assumed six independent lasers, and a simple illustrative example was provided for the case with laser locking. More detailed discussion regarding the impacts and choice of locking schemes also awaits future research.
@@ Appendix A Noise models
Appendix B An illustrative example for locked lasers locking
Figure 8: Similar to FIG. 6, but the green and orange corner plots represent the posteriors of the “no locking” and “N1-LA12 locking scheme” cases, respectively. Both results are obtained based on LFN-only simulations.